OpenStudy (anonymous):

what is the sum of 5/2+(-2/3)

OpenStudy (anonymous):

it is 5/2 - 2/3 =5(3)-2(2)/2(3) =15-4/6 =11/6

OpenStudy (anonymous):

Thanks so much I have struggled with Math all my life

OpenStudy (anonymous):

ur welcome...